OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2).

OpenStudy (anonymous):

y = -one thirdx - 6 y = -one thirdx - 14 y = 3x - 6 y = 3x - 14

OpenStudy (texaschic101):

3x - y = -3 3x + 3 = y (rearrange) y = 3x + 3 the slope of this line is 3 a parallel line will have the same slope now we will use y = mx + b slope(m) = 3 (4,-2)..x = 4 and y = -2 now we will sub and solve for b, the y intercept y = mx + b -2 = 3(4) + b -2 = 12 + b -2 - 12 = b -14 = b so our parallel equation is : y = 3x - 14

OpenStudy (anonymous):

Thank you!!! Can you help me with another one?

OpenStudy (texaschic101):

sure thing :)

OpenStudy (texaschic101):

what ya got ?

OpenStudy (anonymous):

Select the coordinates of two points that are on the line x = 5. (5, 0) and (0, 5) (5, 0) and (5, -2) (5, 0) and (0, 2) (0, 5) and (0, 2)